Consider the following demand and supply equations for sugar:

Qd = 1,000 -1,000P

P = 0.000Qs - 0.8

where P is the price of sugar per pound and Q is thousands of pounds of

sugar.

a. What are the equilibrium price and quantity for sugar? At the same equilibrium price, compute the consumer surplus.

b. Suppose that the government wishes to subsidize sugar production by placing a floor on sugar prices of $0.20 per pound. How much will the government need to spend to ensure that the market clears? Compute the deadweight loss at that point.
1
Expert's answer
2020-10-25T18:24:48-0400

a. The equilibrium price and quantity for sugar are:

Qs = 125P + 100,

Qd = Qs,

1,000 - 1,000P = 125P + 100,

1,125P = 900,

P = 0.8,

Q = 125×0.8 + 100 = 200 units.

The consumer surplus is:

CS = 0.5×(1 - 0.8)×200 = 20.

b. If the government wishes to subsidize sugar production by placing a floor on sugar prices of $0.20 per pound, then it is a non-binding price floor, as the equilibrium price is higher, so there will be no deadwight loss.